# Surgical Gaussian Surfels: Highly Accurate Real-time Surgical Scene Rendering

### v1

@@ Abstract
Accurate geometric reconstruction of deformable tissues in monocular endoscopic video remains a fundamental challenge in robot-assisted minimally invasive surgery. Although recent volumetric and point primitive methods based on neural radiance fields (NeRF) and 3D Gaussian primitives have efficiently rendered surgical scenes, they still struggle with handling artifact-free tool occlusions and preserving fine anatomical details. These limitations stem from unrestricted Gaussian scaling and insufficient surface alignment constraints during reconstruction. To address these issues, we introduce Surgical Gaussian Surfels (SGS), which transforms anisotropic point primitives into surface-aligned elliptical splats by constraining the scale component of the Gaussian covariance matrix along the view-aligned axis. We predict accurate surfel motion fields using a lightweight Multi-Layer Perceptron (MLP) coupled with locality constraints to handle complex tissue deformations. We use homodirectional view-space positional gradients to capture fine image details by splitting Gaussian Surfels in over-reconstructed regions. In addition, we define surface normals as the direction of the steepest density change within each Gaussian surfel primitive, enabling accurate normal estimation without requiring monocular normal priors. We evaluate our method on two in-vivo surgical datasets, where it outperforms current state-of-the-art methods in surface geometry, normal map quality, and rendering efficiency, while remaining competitive in real-time rendering performance. We make our code available at https://github.com/aloma85/SurgicalGaussianSurfels
Keywords:
Gaussian Splatting 3D Surface Reconstruction Minimally Invasive Surgery.
@@ 1 Introduction
Reconstructing dynamic surgical scenes from endoscopic videos is a crucial, yet complex task in robot-aided minimally invasive surgery. Accurate reconstructions improve the precision of surgical instrument manipulation and serve as a foundation for clinical applications such as AR/VR surgical simulation, medical training, and complete medical automation [2]. Factors such as deformable tissues, limited viewing angles, surgical tool occlusion, and severe specularity currently thwart existing 3D surface rendering methods [4].
SGS Result Reference SGS Result Reference
Figure 1: Qualitative results of our method SGS on the StereoMIS [5] Intestine and Liver dataset. We remove surgical tools and render artifact-free surgical tissue.
@@
Advancements in 3D Gaussian Splatting (3DGS) [8] have shown promise in rendering and reconstructing 3D surgical scenes efficiently. Despite these advancements, 3DGS struggles with accurate geometric reconstructions, primarily due to the non-zero thickness of Gaussian point primitives and the ambiguity in normal directions due to their ellipsoid-like shape [1]. To address these limitations, Gaussian Surfels have been proposed, which flatten 3D Gaussian points into 2D ellipses by setting the z-scale to zero [1].
In this paper, we introduce Surgical Gaussian Surfels (SGS), a novel method that combines the surface alignment properties of Gaussian Surfels with the dynamic modeling capabilities of Gaussian Splatting to reconstruct deformable tissues in endoscopic videos. Occluding surgical instruments introduce challenges in reconstruction by obstructing the view of underlying tissues and introducing motion-appearance ambiguities in single-viewpoint scenarios. We address this by incorporating depth priors and a binary motion mask during Gaussian Surfel initialization and training, effectively filtering out occluding instruments and improving reconstruction accuracy.
Our contributions are as follows: (1) We propose SGS, a point-primitive-based rendering technique that takes advantage of the surface-alignment benefits and the high-quality rendering abilities of Gaussian Surfels to achieve high-fidelity reconstruction of dynamic surgical scenes; (2) By defining surface normals as the direction of the steepest density change within each surfel primitive, we accurately predict the orientation of surgical tissue without relying on monocular normal priors, thereby enhancing reconstruction geometry and mitigating shape-radiance ambiguities in highly specular areas; (3) We introduce a Projection-Based Iterative Mask Integration (PIMI) method to initialize point clouds in surgical scenes, resolving gaps and inaccuracies caused by incomplete depth data through confidence-weighted depth aggregation and refined color mapping; (4) SGS outperforms prior implicit and explicit rendering techniques in both quantitative and qualitative experiments.
@@ 2 Related Work
Recent advances in endoscopic 3D reconstruction have been driven by neural-radiance fields and explicit representations. EndoNerf [16] introduces a dual network architecture that combines implicit neural fields with mask-guided ray casting for tool occlusion and depth-cueing ray marching for single-viewpoint reconstruction. EndoSurf [20] advances this framework by representing the canonical scene using signed distance functions (SDF). Neural LerPlane [18] decomposes 4D surgical scenes by factorization into explicit 2D planes of static and dynamic fields to address the computational overhead of volumetric neural fields. EndoGaussian [10] marks a shift towards explicit representations by introducing Holistic Gaussian Initialization (HGI). Endo-4DGS [7] leverages the Depth-Anything foundation model for robust initialization and implements confidence-guided learning to handle uncertain depth measurements in deformable tissues. SurgicalGaussian [17] addresses the limitation of planar structures in encoding complex motion fields by using a deformation network that decouples motion and geometry. Our method builds on and surpasses these prior methods in artifact-free surgical tool removal and tissue reconstruction.
@@ 3 Method
Figure 2: Method pipeline. The full video sequence frames, Binary Mask, Depth Map is passed through PIMI; (a) 2D points are projected to 3D then the tissue surface is represented as Gaussian Surfels with scaling matrix $\mathbf{S}_{k}$, tangent vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and position $\mathbf{x}_{k}$; (b) Deformation prediction MLP takes Surfel position $\mathbf{x}_{k}$ and normalized timestamp $\mathbf{t}$ then predicts the change in position $\delta{x}$, scale $\delta{s}$, and rotation $\delta{q}$; (c) Optimization of Surgical Gaussian Surfels. Pipeline outputs an occlusion-free rendered surgical scene, normal map, and depth map.
@@
Our pipeline, shown in Fig. 2, takes as input a complete monocular endoscopic video sequence, the corresponding binary occlusion masks, and depth maps. The time component $t_{i}$ of each frame $i$ is normalized to $t_{i}=i/T$. We initialize our Surgical Gaussian Surfels from the PIMI (Sec. 3.2) reconstructed point cloud, representing each point as a surface-aligned elliptical splat characterized by a constrained covariance matrix, opacity value, and view-dependent color features represented by spherical harmonic coefficients. To model tissue deformation, we employ a lightweight MLP that predicts per-surfel motion fields while maintaining temporal consistency. The network is set to minimize the photometric differences between the input frames and the rendered views. We introduce several regularization terms to enforce surface smoothness and geometric consistency in occluded regions (Sec. 3.3). Our pipeline enables real-time rendering of dynamic surgical scenes while preserving fine surgical tissue details.
3.1 Preliminaries and Surgical Gaussian Surfels
@@
3D Gaussian Splatting represents a scene as a set of anisotropic 3D Gaussian primitives $\{\mathcal{G}_{k}\mid k=1,\dots,K\}$, each parameterized by its position $\mathbf{p}_{k}\in\mathbb{R}^{3}$, opacity $\alpha_{k}\in[0,1]$, and scaling matrix $\mathbf{S}_{k}\in\mathbb{R}^{3\times 3}$. The covariance matrix of $\mathcal{G}_{k}$ in the world space is defined as $\boldsymbol{\Sigma}_{k}=\mathbf{O}_{k}\operatorname{diag}(\mathbf{s}_{k})\mathbf{O}_{k}^{T}$, where $\mathbf{s}_{k}\in\mathbb{R}^{3}$ is a scaling vector and $\mathbf{O}_{k}\in\mathbb{R}^{3\times 3}$ is a rotation matrix parameterized by a quaternion. The Gaussian function is defined as $\mathcal{G}_{k}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\boldsymbol{\Sigma}_{k}^{-1}(\mathbf{x}-\mathbf{p}_{k})}$. During rendering, Gaussians undergo a view transformation defined by rotation $\mathbf{R}$ and translation $\mathbf{t}$, where $\mathbf{p}_{k}^{\prime}=\mathbf{R}\mathbf{p}_{k}+\mathbf{t},\ \boldsymbol{\Sigma}_{k}^{\prime}=\mathbf{R}\boldsymbol{\Sigma}_{k}\mathbf{R}^{T}$ represent the view-space position and the covariance matrix, respectively. They are then projected to ray space through a local affine transformation $\boldsymbol{\Sigma}_{k}^{2D}=\mathbf{J}_{k}\boldsymbol{\Sigma}_{k}^{\prime}\mathbf{J}_{k}^{T}$, where $\mathbf{J}_{k}$ approximates the projective transformation at the Gaussian center. The view-dependent appearance is modeled using spherical harmonics and alpha composition.
@@
Our Surgical Gaussian Surfels adapt the 3D Gaussian Splatting approach by using surface-aligned elliptical splats, optimized for rendering deformable surgical scenes. Inspired by [6], each surfel $\psi_{k}$ is parameterized by its position $p_{k}\in\mathbb{R}^{3}$, two orthonormal tangential vectors $t_{u},t_{v}\in\mathbb{R}^{3}$, and scaling factors $s_{u},s_{v}\in\mathbb{R}$ that control the size of the Gaussian. The inherent primitive normal is shown in Fig. 2 as $n_{k}=t_{u}\times t_{v}$ represents the direction of the steepest change in surfel density. This formulation enables direct gradient-based optimization, ensuring accurate surfel alignment with the surface of the surgical tissue.
@@
The position of any point on the 2D Gaussian in world space is given by $P(u,v)=p_{k}+s_{u}t_{u}u+s_{v}t_{v}v,\ u,v\in\mathbb{R}$, where $(u,v)$ are local coordinates in the tangent plane. Using homogeneous coordinates, the transformation of the Gaussian from local to world space is represented as:
@@
where $H\in\mathbb{R}^{4\times 4}$ encodes scaling, rotation, zeroed $\mathbf{z}$-component, and translation. As such, $\mathbf{R}_{k}\operatorname{diag}(\mathbf{s}_{k})=H_{k}[:,:3]$. It then follows that the density of each Surgical Gaussian Surfel primitive is modeled as:
@@
Gaussian surfel deformation and rasterization. We use a lightweight MLP to model the deformation field of Gaussian Surfel primitives [17]. The MLP accepts the high-frequency positional and time-encoded [12] surfel primitives corresponding to the current frame as input. The deformation model learns adjustments for various Gaussian properties, such as position, scaling, and rotation, to capture the motion within the scene. Similar to [6], our Surgical Gaussian Surfels are rasterized following a depth-sorted volumetric alpha blending approach:
$c(x)=\sum_{k}c_{k}\alpha_{k}\hat{\mathcal{G}}_{k}(u(x))\prod_{j=1}^{k-1}\left(1-\alpha_{j}\hat{\mathcal{G}}_{j}(u(x))\right),$ (3)
where $c_{k}$ is the color of $k$-th Gaussian, $\alpha_{k}$ its opacity, and $\hat{\mathcal{G}}_{k}$ its adjusted Gaussian value. The process terminates once accumulated opacity reaches saturation.

### v2

@@ Abstract
Accurate geometric reconstruction of deformable tissues in monocular endoscopic video remains a fundamental challenge in robot-assisted minimally invasive surgery. Although recent volumetric and point primitive methods based on neural radiance fields (NeRF) and 3D Gaussian primitives have efficiently rendered surgical scenes, they still struggle with handling artifact-free tool occlusions and preserving fine anatomical details. These limitations stem from unrestricted Gaussian scaling and insufficient surface alignment constraints during reconstruction. To address these issues, we introduce Surgical Gaussian Surfels (SGS), which transform anisotropic point primitives into surface-aligned elliptical splats by constraining the scale component of the Gaussian covariance matrix along the view-aligned axis. We also introduce the Fully Fused Deformation Multilayer Perceptron (FFD-MLP), a lightweight Multi-Layer Perceptron (MLP) that predicts accurate surfel motion fields up to 5× faster than a standard MLP. This is coupled with locality constraints to handle complex tissue deformations. We use homodirectional view-space positional gradients to capture fine image details by splitting Gaussian Surfels in over-reconstructed regions. In addition, we define surface normals as the direction of the steepest density change within each Gaussian surfel primitive, enabling accurate normal estimation without requiring monocular normal priors. We evaluate our method on two in-vivo surgical datasets, where it outperforms current state-of-the-art methods in surface geometry, normal map quality, and rendering efficiency, while remaining competitive in real-time rendering performance. We make our code available at https://github.com/aloma85/SurgicalGaussianSurfels
Figure 1: Left: Our method achieves artifact-free renderings of surgical tool occluded areas on the StereoMIS dataset [9]. Right: 2D plot showing the fast inference speed (fps ↑) and rendering quality (LPIPS ↓) of our method on the StereoMIS [9] Intestine dataset.
@@ 1 INTRODUCTION
Reconstructing dynamic surgical scenes from endoscopic videos is a crucial yet challenging task in robot-aided minimally invasive surgery. Accurate reconstructions not only enhance the precision of surgical instrument manipulation but also serve as a foundation for clinical applications such as AR/VR surgical simulation, medical training, surgical scene analysis, and full medical automation [6]. Factors such as deformable tissues, limited viewing angles, surgical tool occlusion, and severe specularity currently hinder existing 3D surface rendering methods [8].
The volumetric rendering method, NeRF [17], was the first attempt at using a principled ray marching approach to represent surgical scene geometry as a continuous and differentiable volume [27, 30, 12, 16, 24]. However, this method has significant drawbacks, as the surgical scene is implicitly represented as neural network architecture weights and employs a probabilistic ray marching scheme to sample viewing directions, resulting in excessively slow rendering and training times [3].
@@
Advancements in 3D Gaussian Splatting (3DGS) [11] have shown promise in efficiently rendering and reconstructing 3D surgical scenes. By representing scenes with explicit, topology-free Gaussian points, 3DGS facilitates rapid training and real-time rendering, outperforming neural implicit representations in both rendering speed and image quality [4]. Despite these advantages, 3DGS struggles with accurate geometric reconstructions, primarily due to the non-zero thickness of Gaussian point primitives, ambiguity in normal directions due to their ellipsoid-like shape, and difficulties in modeling surface edges [5].
To address these limitations, Gaussian Surfels have been proposed, which flatten 3D Gaussian points into 2D ellipses by setting the z-scale to zero [5]. This transformation resolves the normal ambiguity and aligns the Gaussian points closely with the surfaces of scene objects, significantly improving optimization stability and surface reconstruction fidelity. However, applying Gaussian Surfels directly to dynamic surgical scenes presents new challenges, such as using object-centric anisotropic point primitives to model spatio-temporal deformations of soft tissues and handle surgical instrument occlusions.
In this paper, we introduce Surgical Gaussian Surfels, a novel method that combines the surface alignment capabilities of Gaussian surfels with the dynamic modeling strengths of MLPs for reconstructing deformable tissues in endoscopic videos. Our approach utilizes a fast forward mapping deformation neural network (sec. 3.4) to model the spatiotemporal features of soft tissues, enabling flexible Gaussian motion in observation space. By defining surface normals as the direction of the steepest density change within each Gaussian primitive, we achieve accurate surface normal estimation, thereby improving reconstruction quality and addressing shape-radiance ambiguities in regions with high specular reflections.
As part of our Surgical Gaussian Surfels training pipeline, we incorporate sharp depth maps and boundary masks to enhance geometric supervision and segmentation accuracy. Sharp depths are generated by post-processing raw depth predictions with edge-aware filtering and refinement techniques (Suppl. Sec. 2), resulting in more accurate and artifact-free surface representations. Boundary masks are computed by detecting high-gradient regions in the depth maps, effectively highlighting object contours and transition zones. Integrating these sharp depths and boundary contour masks into our optimization framework provides stronger supervision at object boundaries, leading to improved rendering fidelity and faster training convergence, especially in regions with complex geometry.
To address the unique challenges of surgical environments, we utilize depth priors and a binary motion mask during Gaussian Surfel initialization (sec. 3.3) and training. This effectively removes occluding surgical instruments from the reconstruction and reduces the motion-appearance ambiguity inherent in single-viewpoint scenarios. The efficacy of our method is shown in Fig. 1. Our key contributions are as follows:
  • We propose Surgical Gaussian Surfels, a point-primitive-based rendering technique that combines the object surface alignment properties of surfels [22] with the high-quality explicit rendering quality of 3D Gaussian Splatting to achieve high-fidelity reconstruction of dynamic surgical scenes
  • By defining surface normals as the direction of the steepest density change within each surfel primitive, we accurately predict the orientation of surgical tissue without relying on monocular normal priors, thereby improving reconstruction geometry and mitigating issues caused by specular reflections
  • We propose a Projection-Based Iterative Mask Integration (PIMI) method to initialize point clouds in surgical scenes, which addresses gaps and inaccuracies caused by incomplete depth data by leveraging confidence-weighted depth aggregation and refined color mapping to ensure geometrically consistent and visually accurate 3D reconstructions
  • Based on the tiny-cuda-nn [18] architecture, we develop the Fully Fused Deformation Multilayer Perceptron (FFD-MLP), a high-performance implementation of the MLP that takes advantage of highly parallelized compute architecture to predict surgical tissue deformation fields several times faster than a vanilla MLP
  • In quantitative and qualitative comparisons to previous attempts at dynamic surgical tissue reconstruction, our method outperforms prior implicit and explicit rendering techniques by a wide margin
@@ 2 RELATED WORK
Recent advances in endoscopic 3D reconstruction have been driven by Neural Radiance Fields (NeRFs) and explicit representations. Early neural approaches [27] introduced a dual network architecture that combined implicit neural fields with mask-guided ray casting for tool occlusion and depth-cueing ray marching for single-viewpoint reconstruction. Their method employs Gaussian transfer functions with stereo depth to guide point sampling near tissue surfaces. EndoSurf [33] advanced this framework by representing the canonical scene through signed distance functions (SDFs), employing separate MLPs for deformation, geometry, and appearance. Their geometry network explicitly defines surfaces as zero-level sets, computing surface normals in canonical space and mapping them to observation space through deformation.
Neural LerPlane [30] decomposed 4D surgical scenes by factoring them into explicit 2D planes of static and dynamic fields to address the computational overhead of volumetric neural fields. This approach treats temporal data as 4D volumes with the time axis orthogonal to 3D spatial coordinates, enabling information sharing across timesteps within static fields while reducing the negative impact of limited viewpoints. EndoGaussian [15] marked a shift toward explicit representations by introducing Holistic Gaussian Initialization (HGI) for depth-aware point cloud generation and Spatio-temporal Gaussian Tracking (SGT) for modeling surface dynamics. Their deformation field combines efficient encoding voxels with a lightweight decoder to enable real-time Gaussian tracking.
3DGS-based surgical scene reconstruction methods continue to take advantage of the rapid improvements in the explicit primitives scene reconstruction domain. EndoGS [34] incorporates time-varying deformable fields by representing a surgical scene as a 4D volume. They utilize a multiresolution HexPlane [1, 25] to encode dynamic spatial and temporal information in the surgical scene. Deform3DGS [31] proposes a flexible deformation modeling scheme to learn tissue deformation dynamics for each anisotropic Gaussian primitive. They improve on [28] by adopting learnable mean and variance operators in Fourier and polynomial basis functions to efficiently learn Gaussian point-primitive motion curves.
Figure 2: Method pipeline. The full video sequence frames, Binary Mask, Depth Map is passed through PIMI (sec. 3.3); (a) 2D points are projected to 3D then the tissue surface is represented as Gaussian Surfels with scaling matrix $\mathbf{S}_{k}$, tangent vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and position $\mathbf{x}_{k}$; (b) The Fully Fused Deformation prediction MLP (sec. 3.4) takes Surfel position $\mathbf{x}_{k}$ and normalized timestamp $\mathbf{t}$ then predicts the change in position $\delta{x}$, scale $\delta{s}$, and rotation $\delta{q}$; (c) Optimization of Surgical Gaussian Surfels (sec. 3.5). Pipeline outputs an occlusion-free rendered surgical scene, normal map, and depth map.
Furthermore, SurgicalGaussian [29] addresses the limitation of planar structures in encoding complex motion fields using a deformation network that decouples motion and geometry. Their Gaussian initialization strategy (GIDM) uses geometry priors based on depth and mask information to reduce motion-appearance ambiguity in single-viewpoint scenarios.
Our Surgical Gaussian Surfels method builds on these advances while addressing fundamental geometric accuracy and real-time performance limitations. By constraining the z-scale of anisotropic Gaussian covariance matrices, we preserve fine geometric details that are often lost in traditional 3D Gaussian representations. Our temporal tool masking scheme with depth priors enables robust initialization in canonical space, while our lightweight MLP efficiently maps spatiotemporal deformations. Surgical tissue surface normals are represented as the direction of the steepest density change within each Gaussian primitive, thereby improving surface reconstruction quality while maintaining real-time rendering capabilities.
Figure 3: Qualitative comparison of (a) Our Gaussian surfel distribution-based method (SGS) with correct surface normals, Sharp Depth, and Boundary Mask infused training (Suppl. Sec. 2), (b) a Gaussian Elypsoid-based method SG [29] with ambiguous surface normals.
@@ 3 METHOD
The goal of our method is to achieve high-fidelity geometric reconstruction of deformable surgical scenes by combining the real-time rendering capabilities of 3D Gaussian Splatting with the surface-preserving properties of surfel primitives. As shown in Fig. 2, our pipeline takes as input a temporally aware stack of monocular endoscopic video frames, binary occlusion masks, and depth maps. The time component is normalized to $t_{i}=i/T$. We then initialize our Surgical Gaussian Surfels from the reconstructed point cloud, representing each point as a surface-aligned elliptical splat characterized by a constrained covariance matrix, opacity value, and view-dependent color features represented by spherical harmonic coefficients.
To model tissue deformation, we employ a fast and lightweight MLP that predicts per-surfel motion fields while maintaining temporal consistency. The network is set to minimize photometric differences between input frames and rendered views, without the need for additional monocular normal supervision. We introduce several regularization terms to enforce surface smoothness and geometric consistency in occluded regions (Sec. 3.5). Our pipeline enables real-time rendering of dynamic surgical scenes while preserving fine anatomical details and handling tool-tissue interactions.
3.1 3D Gaussian Splatting
@@
3D Gaussian Splatting represents a scene as a set of anisotropic 3D Gaussian primitives $\{\mathcal{G}_{k}\mid k=1,\dots,K\}$. Each Gaussian $\mathcal{G}_{k}$ is parameterized by its position $\mathbf{p}_{k}\in\mathbb{R}^{3}$, opacity $\alpha_{k}\in[0,1]$, rotation matrix $\mathbf{R}_{k}\in\mathbb{R}^{3\times 3}$, and scaling matrix $\mathbf{S}_{k}\in\mathbb{R}^{3\times 3}$. Together, these form a covariance matrix $\boldsymbol{\Sigma}_{k}=\mathbf{O}_{k}\operatorname{diag}(\mathbf{s}_{k})\mathbf{O}_{k}^{T}$ in world space, where its inverse $\boldsymbol{\Sigma}_{k}^{-1}$ is the precision matrix [14], $\mathbf{s}_{k}\in\mathbb{R}^{3}$ is a scaling vector, and $\mathbf{O}_{k}\in\mathbb{R}^{3\times 3}$ is a rotation matrix parameterized by a quaternion:
$\mathcal{G}_{k}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\boldsymbol{\Sigma}_{k}^{-1}(\mathbf{x}-\mathbf{p}_{k})}$ (1)
For rendering, Gaussians undergo a view transformation defined by rotation $\mathbf{R}$ and translation $\mathbf{t}$:
$\mathbf{p}_{k}^{\prime}=\mathbf{R}\mathbf{p}_{k}+\mathbf{t},\quad\boldsymbol{\Sigma}_{k}^{\prime}=\mathbf{R}\boldsymbol{\Sigma}_{k}\mathbf{R}^{T}$ (2)
They are then projected to ray space through a local affine transformation:
$\boldsymbol{\Sigma}_{k}^{2D}=\mathbf{J}_{k}\boldsymbol{\Sigma}_{k}^{\prime}\mathbf{J}_{k}^{T}$ (3)
where $\mathbf{J}_{k}$ approximates the projective transformation at the Gaussian center. View-dependent appearance is modeled using spherical harmonics, with the final color at each pixel computed through alpha composition:
$c(\mathbf{x})=\sum_{k=1}^{K}\alpha_{k}c_{k}\mathcal{G}_{k}^{2D}(\mathbf{x})\prod_{j<k}\left(1-\alpha_{j}\mathcal{G}_{j}^{2D}(\mathbf{x})\right)$ (4)
3.2 Surgical Gaussian Surfels
3.2.1 Modeling
@@
Our Surgical Gaussian Surfels adapt the 3D Gaussian Splatting approach by introducing surface-aligned elliptical splats, optimized for rendering deformable surgical scenes. Inspired by [5], each surfel $\psi_{k}$ is parameterized by its position $p_{k}\in\mathbb{R}^{3}$, two orthonormal tangential vectors $t_{u},t_{v}\in\mathbb{R}^{3}$, and scaling factors $s_{u},s_{v}\in\mathbb{R}$ that control the size of the Gaussian. The inherent primitive normal is shown in Fig. 2(a) as:
$t_{w}=t_{u}\times t_{v},$ (5)
representing the direction of the steepest change in surfel density. This formulation enables direct gradient-based optimization, ensuring accurate alignment of surfels with the surface of the surgical tissue.
The position of any point on the 2D Gaussian in world space is given by:
$P(u,v)=p_{k}+s_{u}t_{u}u+s_{v}t_{v}v,\quad u,v\in\mathbb{R},$ (6)
@@
where $(u,v)$ are local coordinates in the tangent plane. Using homogeneous coordinates, the transformation of the Gaussian from local to world space is represented as:
@@
where $H\in\mathbb{R}^{4\times 4}$ encodes scaling, rotation, the zeroed $\mathbf{z}$-component, and translation. As such, $\mathbf{R}_{k}\operatorname{diag}(\mathbf{s}_{k})=H_{k}[:,:3]$. It then follows that the density of each Surgical Gaussian Surfel primitive is modeled as: